Status Report

ISS On-Orbit Status 22 Feb 2002

By SpaceRef Editor
February 22, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

Early this morning, the crew reentered the US segment (USOS), after its atmosphere had been cleansed by running the TCCS (trace contaminant control system) all day yesterday. Conditions today returned to normal. Intermodular hatches were opened again, and the IMV (intermodular ventilation) valves were to be opened upon Russian concurrence, predicated on a formal NASA letter with Medical, Safety and Toxicology clearance.

The crew’s evacuation of the USOS to the RS on Wednesday night had been motivated by a foul, sweet smelling odor emanating from the Airlock (A/L) about 2 h 45 min after post-EVA activation of a Metox regeneration oven. Recyclable Metox (metal oxide) canisters are used to remove CO2 and trace contaminants from both the A/L (when closed off from ISS) and the EMU spacesuits. Regeneration of their adsorbent contents requires their heating in a special oven in the A/L, which the crew initiated after ingress. The unpleasant but nontoxic odor came from two canisters that had been installed in the oven for about 7 months. Their regeneration had been initiated shortly after 7A EVA #3, but was aborted 50 min. later to troubleshoot CCAA (common cabin air assembly) noise, and was never re-initiated. Current theory is that inadvertent opening of the oven door exposed the charcoal inside the canisters to trace constituents and allowed it to absorb them for up to 7 months. These contaminants, mostly organic compounds like ethanol, ethyl acetate, isopropanol and toluene, may have been re-released to the cabin atmosphere very quickly, producing the strong odor. In ground testing, MCC-H has reproduced a mild odor similar to the one described by the crew by regenerating a canister that had been sitting un-capped for 10-18 months. A plan is being formulated on how to improve Metox canister recycling in the future.

After a few initial configuration problems, transition to 8A/Release 2 (R2) software on U.S. command and control computers on board has gone generally as planned. By noon today, C&C-3 and C&C-1 MDM were both up and running on the new R2 load (C&C-3 in primary, C&C-1 in backup), and C&C-2, still with R1 s/w, is kept on Standby until tomorrow. GNC-2 MDM (guidance, navigation & control computer #2) is also running nominally on R2, and GNC-1 remains on R1 until tomorrow. Ku-band is back up, and S-band again uses Autotrack for TDRS pointing. As part of the transition, FE-1 Walz swapped out the hard drives of the A/L and Cupola RWS PCS (portable computer systems), installing the new drives with 8A software. Reconfiguring to 8A also involves the POIC (Payload Operations and Integration Center. Huntsville), for whom the transition is a step-up to R3 already. The transition will be completed tomorrow, 2/23.

The crew successfully assisted in correcting a serious power converter configuration mismatch discovered during the 2/20 EVA.Ê [After the ground tried unsuccessfully to command the DDCUs (dc-to-dc converter units), on which EV1 and EV2 had connected cables and closed CIDs (circuit interrupt devices) as planned, they were found to be in the wrong configuration, i.e., set up for later in the assembly sequence (after 12A.1) where they will provide up to 12 kW power to payloads. In that configuration, a parallel control cable allows the DDCUs to operate in a master-slave relationship to provide common voltage regulation and power output. However for 8A, when needed to operate in stand-alone, non-parallel mode, this cable confuses the firmware, causing it to reject all commands.]

The crew this morning rotated the two DDCU racks in the Lab and removed the two offending parallel cables. The ground now will uplink a software patch, telling the DDCUs they are in parallel so they can be commanded back to the stand-alone configuration. An investigation is underway at KSC to determine how this serious configuration mismatch, i.e., DDCU launch configuration significantly different from documentation available to MCC-H, could happen.

FE-2 Bursch completed transferring accumulated condensate (“technical”) water from the Lab tank to a CWC (collapsible water container). Before that, he collected three samples for later analysis, using a special CWC sampling adapter assembly and chemical archive sample bags. Each CWC container holds approximately 45 kg of water (less than its real capacity of 71 kg, to maximize CWC cycle life). The crew was also provided with procedural instructions for torquing the two QDs (quick disconnects) on the AV-1 condensate water vent line’s untorqued “T” connection in the Lab, to prevent them from coming apart during ops. The procedure was completed.

CDR Onufrienko replaced urine collector and filter insert in the ASU toilet system.

Walz and Bursch today completed the important post-EVA PuFF (pulmonary function in flight) experiment session postponed from yesterday due to the USOS evacuation (PuFF was erroneously reported yesterday as complete).

Yesterday, the crew was unable to complete the BINS (Russian strap-down inertial guidance system) correction, using the Puma sighting device and the VP-2 pilot sight, because, according to TsUP/Moscow, the procedure is not fully understood by them. The attempt will be repeated as soon as the crew has reviewed the flight documentation and reports their readiness.Ê [In the meantime‚®Eœumerical BINS updates are accomplished with two other SM sensor systems: three IKVC-256K infrared horizon sensors (which determine orientation by detecting the differential between the warm Earth horizon and cold space), and four SD-251K2 solar sensors. If urgent, the BOKS star sensors could also be used, even though they have caused problems in the past at the current high solar Beta angles. In any case, updating the BINS is currently not considered critical.]

Except for the Metox regeneration, the crew completed reconfiguration of the A/L and EVA equipment stowage.

Dan Bursch completed installation of RPS (rack power switch) decals, for the R2 software load.

Postponed from yesterday due to the contamination issue, the repressurization of the ISS cabin atmosphere to 760 mmHg from the Progress 7P with air (not oxygen, as reported), was accomplished today. The repress from the Progressâ 45-kg high-pressure air tank was done at very slow rate, taking a little over an hour.

Onufrienko and Walz performed an inventory audit using the IMS (inventory management system) that focused mainly on Russian medical support system (SMO) equipment in SM and Lab. On request by Onufrienko, medical specialists had reviewed medical kit shelf lives and uplinked a list of items for disposal. The crew was also requested to provide IMS updates for some undocumented items such as a harmful contaminant filter (FVP), a gas mask (MPK) and free volume availability.

Carl Walz completed reading and downloading radiation dose data from the twelve EVARM (EVA radiation monitoring) badges that were collected Wednesday outside the station. The data were then transferred from the EVARM Reader to the HRF (human research facility) PC hard drive for later downlinking.

During the EVA, EV2 noted that one of his gloves had been contaminated with a substance from a tether, installed several flights ago. The two crewmembers had no wipes with them; when back inside the A/L, they wiped the glove and sealed glove and wipe in a bag. The substance is believed to be lubricant or flaking due to oxidation of the paint on the reel. As a precaution, the glove will no longer be used and will be returned on a future Shuttle flight. The EVA Office will also implement a standard requirement for spacewalkers to have access to wipes during U.S. EVAs.

RSC-Energia reports that launch of Progress M1-257 (7P) is now officially set for 3/21 (backup date: 3/22). Undocking of 6P from ISS is scheduled before the launch, on 3/19, leaving sufficient time for conducting the Kolibri microsatellite experiment prior to 6P atmospheric entry. Trash bags containing soft/liquid trash were found to be safe when the Progress cargo compartment is depressurized to vacuum using the regular pressure relief valve. Depress will take place prior to undocking.Ê [Original concept for controlling two Progress vehicles was to undock first, then launch the second. Later, to protect the Mir docking port from overheating during unprotected sun exposure, the concept was changed to launch first, then undock. Now, with no thermal constraint on the SM aft port, TsUP returns to the first concept. Primary constraint is phasing: to prevent possible interference of radio signals going to different Progress vehicles, launch is OK if the central phasing angle is less than 300 deg (i.e., the second, orbiting Progress below the horizon of the launch site). When the angle is close to 360 deg (i.e., equal to 0 deg), launch is to be avoided. If launch of 7P is delayed, 6P will be kept on the station to retain the two-day interval between launch and undocking.]

Today’s CEO targets: Lake Nasser, Toshka Lakes, Egypt (an excellent pass with near-nadir views for mapping the Toshka Lakes. These man-made are filling with water diverted from the Nile River. Of interest: documenting the water color and shoreline changes of this expanding series of lakes west of the Nile), Sahel Fires, West Africa (Dynamic Event Site: After nearly a decade of moist years in western Africa, last year has been a dry one and fires are becoming widespread in the dry vegetation. Crew to use 50 and 110 mm lens looking obliquely to either side of track to detect the location and extent of fires and their smoke plumes), W. Mediterranean Dust and Smog (weak high pressure with light winds is holding on over the western Mediterranean. Of interest: oblique views either side of track detect the presence and build up of atmospheric aerosols in this region), Eastern United States (as high pressure builds into the southeastern US, crew was to point their camera to the left of track for smoke and smog build up over the Piedmont Plateau and the coastal waters between northern Florida and Cape Hatteras), and Lakes of the Eastern Sierra Watershed (crew was advised to look to the left of track up the spine of the Sierra Nevadas and the desert areas to their east for context photos of the snow pack and playas or water bodies visible there).

U.S. and Russian Segment Status (as of 2:05 pm EST today):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On (16 Amp mode, the lowest possible setting).
  • Vozdukh CO2 scrubber is in MANUAL cycle mode #5 (vacuum pump failed). U.S. CDRA CO2 scrubber is in Standby.

  • BMP Harmful Impurities unit: Both absorbent beds (Filters #1 & #2) in Purify mode.

  • SM Working Compartment: Pressure (mmHg) — 764, temperature (deg C) — 26.0, ppO2 (mmHg) — 160.9, ppCO2 (mmHg) — 2.5 (? see Notes below and above)

  • SM Transfer Compartment: Pressure (mmHg) — 762, temperature (deg C) 21.2; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

  • FGB Cabin: Pressure (mmHg) — 756, temperature (deg C) — 21.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

  • Node: Pressure (mmHg) — 760.43, temperature (deg C) — 23.3 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

  • U.S. Lab: Pressure (mmHg) — 762.19, temperature (deg C) — 22.3, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;

  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 761.89, temperature (deg C) — 21.8; shell heater temp (deg C) — 21.9, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

  • PMA-1: Shell heater temp (deg C) — 23.3.

  • PMA-2: Shell heater temp (deg C) — 12.3.

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment (USOS) not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]. Russian GA (gas analyzer) readings of ppCO2 in the SM are invalid).

Electrical Power Systems (EPS):

  • Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in Autotrack mode (during XPOP).

  • SM batteries: all batteries (8) in “partial charge” mode.

  • FGB battery #5 is offline; battery #4 is cycling; all other (4) batteries are in “partial charge” mode.

  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is Off (Freon leak). SKV-2 is On.

Command & Data Handling Systems:

  • C&C-3 MDM is prime, C&C-1 is back-up, and C&C-2 is in standby, with R1 s/w.

  • GNC-2 MDM is prime; GNC-1 is back-up, with R1 s/w.

  • LA-1, LA-2 and LA-3 MDMs are all operating.

  • APS-1 (automated payload switch #1) and APS-2 are both On.

  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.

  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:

  • S-band is operating nominally.

  • Ku-band is operating nominally in open loop pointing mode.

  • Audio subsystem operating nominally.

  • Video subsystem operating nominally.

  • MCOR (medium-rate communications outage recorder) operating nominally.


  • SSRMS/Canadarm2 at Port Stow position, with Keep Alive power on both strings.

  • RWS (robotics workstations) are Off.

SpaceRef staff editor.